What is 5(3+4)/35+12/4
7+{20-(7+6)}
9+{13(2)-(3-12)}
12/2-{6-(7(3))}

Answer by checkley77(12844)

5(3+4)/35+12/4
5*7/35=3
35/35+3
1+3=4 ANS.
-----------------
7+{20-(7+6)}
7+20-13
27-13=14 ANS.
-----------------
9+{13(2)-(3-12)}
9+{26-(-9)}
9+{26+9}
9+35=44 ANS.
-------------------
12/2-{6-(7(3))}
6-{6-21}
6-{-15}
6+15=21 ANS.